OpenStudy (anonymous):

Solve for x: HELP! x = –4 x = 5 x = –4 and x = 5 No Solution

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }+\frac{ 4 }{2x }=\frac{ x+4 }{ 10 }\]

OpenStudy (anonymous):

make 1/2 have a denominator of x, add the two terms on the left hand side, then cross multiply

OpenStudy (anonymous):

sorry, a denominator of 2x

OpenStudy (anonymous):

add both 2xs?

OpenStudy (anonymous):

well, if you multiply 1/2 by x/x, you get x/2x. You can then add x/2x and 4/2x

OpenStudy (anonymous):

4x/2x? or 4x/4x?

OpenStudy (anonymous):

the first one now cross multiply

OpenStudy (anonymous):

actually no, it's 4 + x my bad

OpenStudy (anonymous):

that is, 4+x/2x

OpenStudy (anonymous):

what do i cross multiply?

OpenStudy (anonymous):

\[\frac{4+x}{2x} = \frac{x+4}{10}\]

OpenStudy (anonymous):

x=-4